OpenStudy (anonymous):

what is the 5th term of an arithmetic sequence if

OpenStudy (anonymous):

-.- t2=-5 and t6=7

OpenStudy (anonymous):

5, 3, 6, 4 I don't understand how I can find the first term to this.

OpenStudy (anonymous):

like don't we need the first term to solve it?

OpenStudy (anonymous):

We have two ways to solve this exercise, I'll chose one way: \[a_n=a_x+(n-x)*r\]\[a_6=a_2+(6-2)*r\]\[7=-5+4*r\]\[r=3\]So\[a_1=a_2+(1-2)*3\]\[a_1=-5+(-1)*3\]\[\boxed{a_1=-8}\]

OpenStudy (anonymous):

not understanding what you're saying.

OpenStudy (anonymous):

oooh jk. I think I'm starting to understand it.

OpenStudy (anonymous):

okay okay so I got to -5=t1+1d 7=t1+5d you have to subtract those right?

OpenStudy (anonymous):

The other way is:\[a_n=a_1+(n-1)r\] \[a_2=a_1+r\\a_6=a_1+5r\]then you solve this system for r and a1